Epidemiology question: In population C a total of 4,350 people were diagnosed with influenza during flu season (roughly 6 months). Of patients with the flu, 27 of them died while they had the flu; however one of these fatalities was due to a car accident.

2a. What was the case fatality rate? Report as per 1000. (1pt.)
Answer #1

total number of people diagnosed with influenza = 4350

number of people died = 27 - 1 = 26 died due to flu

Case fatality rate = [(number of people died)/(total number of people diagnosed with influenza)]*1000

= (26/4350)*1000

= 5.98 per 1000 is required case fatality rate (2 decicmals)